A professor at a college analyzed the relationship between the final grade in Calculus and factors affecting its achievement with a sample of 80 students. The independent variables included in the regression model are as follows: x1: Final grade for College Algebra, x2: ACT math score, x3: ACT natural science score, xs: Percentile high school rank. The following ANOVA summarizes the regression results. Table 1: ANOVA Source of Variation df Regression Residual or Error Total Source of Squares Mean Square F 2986.150 7,346.050 10,332.200 Question 1: (1 points) a) What is the degrees of freedom for Regression? b) What is the degrees of freedom for Residual? Question 2: (1 points) What is the value of the Regression mean square (MSR)? Round the answer to the nearest 1000th MSR- What is the value of the Error mean square (MSE)? Round the answer to the nearest a) b) 1000th MSE-
Answer #1

1)

a) degree of freedom for regression =number of independent variables p=4

b) degree of freedom for residual =n-p=80-4-1 =75

2)

a) MSR =SSR/df(regression)=2986.150/4=746.538

b) MSE =7346.05/75=97.947